# Ultra-Efficient Decoding for End-to-End Neural Compression and Reconstruction

### v1

@@ 3 Framework for Ultra-Efficient Reconstruction
Singular-Value Decomposition (SVD) decomposes a matrix AA into three matrices, $\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{T}$, where $\mathbf{U}$ and $\mathbf{V}$ are orthogonal matrices, and $\mathbf{\Sigma}$ is a diagonal matrix of singular values Beltrami1873 (6). A low-rank approximation of any matrix $\mathbf{A}$ can be obtained by retaining only the top $R$ singular values and their corresponding singular vectors, as shown by the Eckart-Young theorem Eckart_Young_1936 (7). This principal can be extended wherein a matrix is iteratively approximated by observing the residual error among a series of summed low-rank reconstructions.
@@
where $\mathbf{U}_{i}\in\mathbb{R}^{m\times m},\mathbf{\Sigma}_{i}\in\mathbb{R}^{m\times n},\mathbf{V}^{T}_{i}\in\mathbb{R}^{n\times n}$ and $\mathbf{U}_{i_{R}}\in\mathbb{R}^{m\times R},\mathbf{\Sigma}_{i_{R}}\in\mathbb{R}^{R\times R},\mathbf{V}^{T}_{i_{R}}\in\mathbb{R}^{R\times n}$, with $R$ corresponding to the target rank $R\in[1,\text{min}(m,n)]$, and $\mathbf{A}_{i_{R}}$ denoting the low rank version of any matrix $\mathbf{A}$. Following Eq. 3, once all II low-rank approximations have been made, they produce the final reconstruction $\hat{\mathbf{T}}=\frac{1}{I}\sum_{i=0}^{I}\mathbf{T}_{i_{R}}$. The reconstruction can be evaluated by comparing $\hat{T}$ with $T_{0}$, where the element-averaged p-norm error is given by $\frac{1}{N}(\sum_{elems}(|\mathbf{T}_{0}-\hat{\mathbf{T}}|^{p}))^{\frac{1}{p}}$.

### v2

@@ 3 Framework for Ultra-Efficient Reconstruction
Singular-Value Decomposition (SVD) decomposes a matrix $A$ into three matrices, $\mathbf{U}\mathbf{\Sigma}\mathbf{V}^{T}$, where $\mathbf{U}$ and $\mathbf{V}$ are orthogonal matrices, and $\mathbf{\Sigma}$ is a diagonal matrix of singular values Beltrami1873 (6). A low-rank approximation of any matrix $\mathbf{A}$ can be obtained by retaining only the top $R$ singular values and their corresponding singular vectors, as shown by the Eckart-Young theorem Eckart_Young_1936 (7). This principal can be extended wherein a matrix is iteratively approximated by observing the residual error among a series of summed low-rank reconstructions.
@@
where $\mathbf{U}_{i}\in\mathbb{R}^{m\times m},\mathbf{\Sigma}_{i}\in\mathbb{R}^{m\times n},\mathbf{V}^{T}_{i}\in\mathbb{R}^{n\times n}$ and $\mathbf{U}_{i_{R}}\in\mathbb{R}^{m\times R},\mathbf{\Sigma}_{i_{R}}\in\mathbb{R}^{R\times R},\mathbf{V}^{T}_{i_{R}}\in\mathbb{R}^{R\times n}$, with $R$ corresponding to the target rank $R\in[1,\text{min}(m,n)]$, and $\mathbf{A}_{i_{R}}$ denoting the low rank version of any matrix $\mathbf{A}$. Following Eq. 3, once all $I$ low-rank approximations have been made, they produce the final reconstruction $\hat{\mathbf{T}}=\frac{1}{I}\sum_{i=0}^{I}\mathbf{T}_{i_{R}}$. The reconstruction can be evaluated by comparing $\hat{T}$ with $T_{0}$, where the element-averaged p-norm error is given by $\frac{1}{N}(\sum_{elems}(|\mathbf{T}_{0}-\hat{\mathbf{T}}|^{p}))^{\frac{1}{p}}$.